# ACT-ing on inflation: Implications of non Bunch-Davies initial condition and reheating on single-field slow roll models

### v1

@@ VI Conclusions
For chaotic inflation, we observe that with a low reheating equation of state parameter $w$, monomial potentials of the form $\phi^{1/5}$ and $\phi^{1/3}$ can satisfy the $1\sigma$ bounds on $n_{{}_{\mathrm{S}}}$ and $r$, while the $\phi^{1/2}$ potential lies within the $2\sigma$ region. However, the canonical slow roll scenario with a quadratic potential $m^{2}\phi^{2}$ remains strongly disfavored by the data. Upon introducing a NBD initial state, we find that the required value of $\beta_{0}$ to achieve $n_{{}_{\mathrm{S}}}=0.9743$ can be either significantly positive or negative, depending on the model parameters. Despite this, the incorporation of NBD initial conditions does not substantially improve the compatibility of chaotic inflation with observations. In particular, for the $\phi^{1/3}$ potential, which was previously consistent with the data once reheating effects were included, the NBD correction tends to increase $r$ beyond the observational upper limit. This suggests that in order to reconcile chaotic inflation with current constraints, one may need to consider extensions involving non-canonical kinetic terms or modified gravity frameworks.
@@
Lastly, natural inflation remains in significant tension with the current observational data, even after incorporating a phase of reheating or introducing a NBD initial condition. A similar conclusion holds for chaotic inflation with higher powers of the potential, such as the quadratic case. These findings indicate that modest extensions within the canonical framework are insufficient to reconcile these models with the data. Therefore, more substantial modifications, such as introducing non-canonical kinetic terms, exploring multi-field inflation, may be necessary for these classes of inflationary models to remain observationally viable.
@@ Acknowledgments
SM thanks the Indian Institute of Science Education and Research (IISER), Pune for support through a postdoctoral fellowship. SM also wishes to thank Arka Banerjee, Diptimoy Ghosh, Susmita Adhikari, Sachin Jain and Farman Ullah for the useful discussions.

### v2

@@ VII Conclusions
For chaotic inflation, we observe that with a low reheating equation of state parameter $w_{\mathrm{re}}$, monomial potentials of the form $\phi^{1/5}$ and $\phi^{1/3}$ can satisfy the $1\sigma$ bounds on $n_{{}_{\mathrm{S}}}$ and $r$, while the $\phi^{1/2}$ potential lies within the $2\sigma$ region. However, the canonical slow roll scenario with a quadratic potential $m^{2}\phi^{2}$ remains strongly disfavored by the data. Upon introducing a NBD initial state, we find that the required value of $\beta_{0}$ to achieve $n_{{}_{\mathrm{S}}}=0.9743$ can be either significantly positive or negative, depending on the model parameters. Despite this, the incorporation of NBD initial conditions does not substantially improve the compatibility of chaotic inflation with observations. In particular, for the $\phi^{1/3}$ potential, which was previously consistent with the data once reheating effects were included, the NBD correction tends to increase $r$ beyond the observational upper limit. This suggests that in order to reconcile chaotic inflation with current constraints, one may need to consider extensions involving non-canonical kinetic terms or modified gravity frameworks.
@@
Lastly, natural inflation remains in significant tension with the current observational data, even after incorporating a phase of reheating or introducing a NBD initial condition. A similar conclusion holds for chaotic inflation with higher powers of the potential, such as the quadratic case. These findings indicate that modest extensions using NBD initial state within the canonical framework are insufficient to reconcile these models with the data. Therefore, more substantial modifications, such as introducing non-canonical kinetic terms, exploring multi-field inflation, may be necessary for these classes of inflationary models to remain observationally viable.
@@ Acknowledgments
SM thanks the Indian Institute of Science Education and Research (IISER), Pune for support through a postdoctoral fellowship. SM also wishes to thank Arka Banerjee, Diptimoy Ghosh, Susmita Adhikari, Sachin Jain, and Farman Ullah for useful discussions. SM would also like to thank Shahin Sheikh Jabbari, Amjad Ashoorioon, and H.V. Ragavendra for their valuable comments.